Ellen is using different colored rectangular cards for her project. Each card has an area of 24 square inches and the length is 10 inches more than the width. Find the length and width of her cards.
If x represents the width of each card, select all the equations that could be used to solve this problem.

(x+10)(x)=24

2(x)+2(x+10)=24

x(10x)=24

x(x+10)=24
Part 2
Find the dimensions of each rectangular card.
width = inches
length = inches

Answer by stanbon(75887)

Ellen is using different colored rectangular cards for her project. Each card has an area of 24 square inches and the length is 10 inches more than the width. Find the length and width of her cards.
If x represents the width of each card, select all the equations that could be used to solve this problem.
Answers::
(x+10)(x)=24
x(x+10)=24
------------------------
Part 2
Find the dimensions of each rectangular card.
x(x+10) = 24
x^2 + 10x - 24 = 0
(x+12)(x-2) = 0
------
width = 2 inches
length = 12 inches
------
Cheers,
Stan H.
-----------